OpenStudy (anonymous):

Compute The following derivatives. Please explain step by step (x^3+e^2x)'=

OpenStudy (skyz):

@mathslover

hartnn (hartnn):

know the derivative of x^n ?

OpenStudy (anonymous):

can you show me step by step @hartnn

hartnn (hartnn):

First we split the derivative, \(\Large (x^3 + e^{2x})' = (x^3)' + (e^{2x})'\)

hartnn (hartnn):

then since we know \((x^n)' = n x^{n-1}\) \(\Large (x^3)' = 3x^{3-1} = 3x^2\)

hartnn (hartnn):

and \((e^x)'=e^x\) but since we have 2x as exponent, we need to use chain rule. \(\large (e^{2x})' = e^{2x} (2x)' = 2e^{2x}\) so, your derivative would finally be \(\Large 3x^22e^{2x}\)

hartnn (hartnn):

i mean \(\Large 3x^2+2e^{2x}\)

OpenStudy (shadowlegendx):

Good job @hartnn

OpenStudy (anonymous):

ok can you tell me why do you split the two above 1ST?

OpenStudy (shadowlegendx):

50_cent msg hartnn so he can continue to assist you...I'm sorry but I don't know this stuff >.>

hartnn (hartnn):

because we can split it up, \((A+B)' = A'+B'\) its the rule in differentiation. sum of differentials, is the differentials of sum

OpenStudy (anonymous):

can you help me to understand how to use d/dt ((5-2t)(t+1))?

hartnn (hartnn):

multiply them out first (5-2t)(t+1) foil

OpenStudy (isaiah.feynman):

Or still use the product rule.

OpenStudy (isaiah.feynman):